Two trains leave stations
156
miles apart at the same time and travel toward each other. One train travels at
55
miles per hour while the other travels at
75
miles per hour. How long will it take for the two trains to meet?

Answer by Edwin McCravy(20081)

Their approach rate is the sum of their speeds, or 55+75 = 130 mph.

The distance between them shrinks from 156 miles to 0 miles at that rate.

d = rt

t = d/r

t = 156/130 = 1.2 hours

Edwin